A 100cm3 gas cylinder filled with chlorine under 160 torr pressure is connected by stop corck with another cylinder of 400cm3 filled nitrogen under pressure of 200torr .what will be the total pressure when stop corck is opened?
1
Expert's answer
2020-11-09T14:06:41-0500

Final volume = (100+400)cm3 = 500cm3

For chlorine, P1V1 =P2 V2

160×100= P2 ×500

P2 = 160×100/500 = 32 torr


For nitrogen, P1V1 =P2 V2

400×200= P2 ×500

P2 =400×200/500= 160 torr


Total pressure = (32+160) torr = 192 torr